# Chain-of-Query: Unleashing the Power of LLMs in SQL-Aided Table Understanding via Multi-Agent Collaboration

### v1

@@ Abstract
Table understanding requires structured, multi-step reasoning. Large Language Models (LLMs) struggle with it due to the structural complexity of tabular data. Recently, multi-agent frameworks for SQL generation have shown promise in tackling the challenges of understanding tabular data, but existing approaches often suffer from limitations such as the inability to comprehend table structure for reliable SQL generation, error propagation that results in invalid queries, and over-reliance on execution correctness. To address these issues, we propose Chain-of-Query (CoQ), a novel multi-agent framework for SQL-aided table understanding. CoQ adopts natural-language-style representations of table schemas to abstract away structural noise and enhance understanding. It employs a clause-by-clause SQL generation strategy to improve query quality and introduces a hybrid reasoning division that separates SQL-based mechanical reasoning from LLM-based logical inference, thereby reducing reliance on execution outcomes. Experiments with four models (both closed- and open-source) across five widely used benchmarks show that Chain-of-Query significantly improves accuracy from 61.11% to 74.77% and reduces the invalid SQL rate from 9.48% to 3.34%, demonstrating its superior effectiveness in table understanding. The code is available at https://github.com/SongyuanSui/ChainofQuery.
@@
Songyuan Sui1, Hongyi Liu1, Serena Liu1, Li Li2, Soo-Hyun Choi2, Rui Chen2, Xia Hu1 1Rice University {Songyuan.Sui, Hongyi.Liu, Serena.Liu, Xia.Hu}@rice.edu 2Samsung Electronics America {li.li1, sh9.choi, rui.chen1}@samsung.com
@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable performance across a variety of natural language processing (NLP) tasks (Yang et al., 2023). However, they still struggle with understanding tabular data (Sui et al., 2023a). This challenge stems from two main factors. First, the structure of tabular data differs significantly from that of plain text. It organizes information via rows and columns, introducing hierarchical relationships, positional dependencies, and implicit semantics that are challenging for language models to capture. Second, table-based tasks often involve multi-step, structured reasoning operations, such as aggregation, comparison, and arithmetic computation (Lu et al., 2025). They go beyond surface-level language understanding.
To adapt LLMs for table understanding, prior work has explored many approaches. Many studies (Herzig et al., 2020; Zhang et al., 2023a, b; Li et al., 2024b; He et al., 2025) fine-tuned LLMs on table-specific datasets, but these methods treat the task as a single-turn generation problem, which results in shallow reasoning paths and overlooks intermediate data. Recent studies (Wang et al., 2024c; Ji et al., 2024; Zhou et al., 2025) built agent-based pipelines with table manipulation tools to conduct multi-step reasoning. Nevertheless, these tools rely on human-defined programs whose functionalities are restricted and cannot perform complex or adaptive operations beyond their predefined scope.
@@
We argue that incorporating Structured Query Language (SQL) provides a more principled approach for understanding tabular data. SQL is inherently designed to access, filter, and aggregate information from relational tables, making it well-aligned with the needs of table understanding. Although existing work on agent-based SQL generation has made some progress in table understanding (Ye et al., 2023; Kong et al., 2024a), its performance remains suboptimal. We identify three key challenges in SQL-aided table understanding with LLMs. 1. Existing approaches typically feed the table and task instructions into an LLM, assuming the LLM can understand the table in order to produce meaningful SQL. However, if the LLM already struggles with table understanding, it is unlikely to generate accurate SQL in the first place. This workflow thus inherits LLMs’ structural inability, as effective SQL generation still requires a solid grasp of the table’s layout and semantics—something LLMs inherently lack. 2. Supporting complex, multi-step reasoning often requires SQL queries with multiple clauses and nested structures. Such queries are difficult to construct correctly in a single shot, and even small errors can cascade into query failure. 3. Existing methods depend heavily on the correctness of SQL execution, as they directly use SQL results as final answers. This further leads LLMs to generate overly complicated queries, reducing reliability.
@@
To tackle these challenges, we propose Chain-of-Query (CoQ), a novel multi-agent framework for SQL-aided table understanding. To address the first challenge, CoQ replaces raw table inputs with natural-language representations of table schemas. Our insight is that SQL generation depends primarily on high-level schema, which can be easily and accurately expressed in natural language, bypassing the need for LLMs to interpret complex tabular structures. This abstraction allows LLMs to focus on semantics without being hindered by noisy or irregular layouts. To address the second challenge, we propose a Clause-by-Clause SQL Generation Strategy, which incrementally constructs SQL queries one clause at a time, forming a chain of progressively refined queries. To solve the third challenge, we design a Hybrid Reasoning Division Strategy that separates mechanical reasoning (executed via SQL) from logical reasoning (delegated to LLMs), treating SQL outputs as intermediate steps rather than final answers. In addition, our Parallel Decomposition enhances robustness by splitting complex questions into parallelizable sub-questions, avoiding the inter-step dependencies of traditional sequential reasoning. Together, CoQ offers a novel design that directly addresses these challenges, establishing an effective new paradigm for table understanding.
@@
To comprehensively evaluate the effectiveness of CoQ, we conduct extensive experiments using both the closed-source models (GPT-3.5, GPT-4.1) and the open-source models (LLaMA 2, DeepSeek-V3), across five representative table understanding benchmarks: WikiTQ, TabFact, FeTaQA, IM-TQA, and Open-WikiTable. Experimental results show that CoQ consistently outperforms state-of-the-art (SOTA) baselines, increasing accuracy from 61.11% to 74.77%, and reducing the invalid SQL generation rate from 9.48% to 3.34%.
@@
  • Extensive experiments show Chain-of-Query consistently outperforms previous baselines, achieving substantial improvements across diverse settings.
@@ 2 Related Work
We review previous work from three aspects: Traditional LLM-Based Table Understanding, Multi-Agent Table Understanding, and Text-to-SQL.
Traditional LLM-Based Table Understanding. Since the structure of tables (e.g., non-sequential cell order) differs significantly from plain text typically used in pre-training (Raffel et al., 2023), some work (Zhang et al., 2023a, b; Li et al., 2024b; Zhuang et al., 2024; He et al., 2025) fine-tuned LLMs to adapt to tabular data. While effective on specific datasets, these methods depend on crafted instructions and expensive training, limiting scalability. More recent prompt-based approaches (Sui et al., 2023a, b; Chen et al., 2024) exploit LLMs’ general reasoning ability, offering better generalization across tasks. However, these methods typically adopt single-turn prompting strategies, failing to support multi-hop reasoning.
Multi-Agent Table Understanding. Recent work has explored LLM-powered agents (Wu et al., 2023; Wang et al., 2024a) that perform multi-step reasoning and invoke table operation tools. Dater (Ye et al., 2023) reformulates questions into cloze-style prompts and retrieves values via LLM-generated SQL. Chain-of-Table (Wang et al., 2024c) and Tree-of-Table (Ji et al., 2024) construct dynamic reasoning paths using pre-defined Python functions. Although effective, these approaches rely on coding-based steps. This makes them fragile, as code errors can cause the entire reasoning chain to fail. For example, Zhou et al. (2025) reports that about half of the failures come from invalid code generation. Moreover, most of these methods are limited by fixed function sets, restricting their reasoning flexibility.
Text-to-SQL. Generating SQL queries from natural language is a long-standing challenge. Early methods primarily rely on fine-tuning LLMs on annotated SQL datasets (Pourreza and Rafiei, 2024; Li et al., 2024a). To reduce training cost and improve compositional reasoning, Tai et al. (2023) and Pourreza and Rafiei (2023) explored Chain-of-Thought prompting in Text-to-SQL. OpenTab (Kong et al., 2024a) integrates LLM-based SQL generation into a tabular data RAG system. Recently, MAC-SQL (Wang et al., 2025b) and MAG-SQL (Xie et al., 2024b) build multi-agent frameworks that refine generated queries to improve SQL quality. Although these methods perform well on benchmarks such as Spider (Yu et al., 2019) and BIRD (Li et al., 2023), they all tend to generate fully grounded SQL queries that directly return the final answer, making query construction difficult. Moreover, LLMs tend to produce overly complex queries in an attempt to handle questions holistically (Shen et al., 2025), reducing reliability.
@@ 3.3 Semantic Splitter: Schema Abstraction and Query Decomposition
The Semantic Splitter serves as the entry point of the entire agentic workflow. This module is designed to optimize the input for SQL generation by addressing two key challenges: (1) LLMs’ limited ability to interpret complex table structures, which often leads to inaccurate SQL generation; and (2) interference between sub-questions within a single complex query, where independent sub-questions could be handled separately to avoid entanglement. To address these challenges, we introduce a natural-language-style representation of table schemas for query generation, and incorporate a parallel decomposition mechanism to identify and isolate separable sub-queries for clean execution. The following subsections detail this process.
@@ 3.3.1 Natural-Language-Style Table Schema
With the goal of addressing these limitations in handling raw and irregular table structures, we construct a concise natural-language-style schema as a surrogate for full-table input. It consists of three components: table name, headers, and value examples. The table name employs concise keywords to summarize the main content and purpose of the table at a high level, providing essential context for the LLM to understand the table. The table headers comprise multiple (column name, column value type) pairs. These pairs not only provide a natural language abstraction of the table structure, but also serve as a reference for value formatting in downstream SQL generation. The value examples are a few sampled rows of the table paired with their corresponding column names. We adopt "PIPE" format (i.e., a pipe-delimited string with "|" separators) to encode the value examples, as studies found it to improve LLM performance on table tasks (Sui et al., 2023a; Wang et al., 2024c).
By replacing the full table with our designed schema, LLMs can generate accurate SQL without being distracted by complex layouts or unnecessary content. This approach also improves scalability, as it avoids importing long tables in full, reducing input length and simplifying the reasoning process.
@@ 3.3.2 Parallel Decomposition
The second component of the Semantic Splitter is a Parallel Decomposer, which focuses on separating input questions into independent sub-questions. Unlike the traditional sequential decomposition strategies where each sub-question depends on the previous answer and introduces fragility due to error propagation, our approach targets questions without interdependencies. We defer the handling of sequential dependencies to our Clause-by-Clause Generator in Section 3.4.2.
Many complex tabular questions involve comparing or aggregating information from semantically distinct and disjoint table regions, which do not require sequential reasoning. To address this, the Parallel Decomposer splits such questions into independent sub-questions, each focused on a localized region of the table. This removes inter-step dependencies, enables concurrent processing, and simplifies SQL generation, as each sub-query operates over a narrow, self-contained scope without interference from others.
@@ 3.4 Clause-wise SQL Generation for Precision and Robustness
After outlining our approach to transforming and preprocessing inputs to improve LLMs’ understanding, we now focus on the next key component: generating accurate SQL with LLMs to support effective table understanding. The SQL Query Generator constructs executable SQL queries to retrieve intermediate information relevant to the task and support mechanical reasoning. As shown in Figure 2, an initial query is generated by an LLM, then validated before being passed to the Dynamic Planner, which determines whether further clauses should be appended to improve precision. We now detail the generation process.
@@ 3.4.1 SQL Query Generation with Validation
Given the schema and sub-question, an LLM generates an SQL query to retrieve relevant information. The query is validated by an SQL executor through direct execution; if it fails, the error message is passed to a secondary LLM for correction. The revised query is re-executed and, if successful, forwarded to the Dynamic Planner. To enhance query quality, we introduce a clause-by-clause generation strategy described below.
@@ 3.4.2 Clause-by-Clause SQL Generation
To improve the reliability and controllability of SQL generation, we propose a Clause-by-Clause Strategy that incrementally builds queries by isolating the construction of each clause. The generator starts by producing a basic SELECT-FROM query that selects columns relevant to the sub-question. However, without condition filtering, this query may return a large number of rows, posing scalability challenges. To mitigate this, we sample a few representative rows (similar to the natural-language-style table schema construction) as the Planner only needs to observe data patterns rather than precise values.
Subsequent clauses (e.g., WHERE, GROUP BY) are appended one at a time, under the guidance of the Dynamic Planner. By decomposing SQL construction into discrete clauses, each step focuses solely on its own logic and syntax, independent of the full query context. Each intermediate query is validated before proceeding. If execution fails, the error can be traced to the most recently added clause, allowing for more targeted correction. If correction fails, the agent reverts to the last validated query, ensuring robustness. As shown in Figure 2, if a newly generated clause introduces an error to the query (SQL2) and cannot be resolved, the agent uses the last valid query (SQL1) to ensure executable SQL. This incremental process forms a chain of increasingly precise queries, where each step builds on a validated state, preventing error propagation and ensuring stability.

### v3

@@ Abstract
Table understanding requires structured, multi-step reasoning. Large Language Models (LLMs) struggle with it due to the structural complexity of tabular data. Recently, multi-agent frameworks for SQL generation have shown promise in tackling the challenges of understanding tabular data, but existing approaches often suffer from limitations such as the inability to comprehend table structure for reliable SQL generation, error propagation that results in invalid queries, and over-reliance on execution correctness. To address these issues, we propose Chain-of-Query (CoQ), a novel multi-agent framework for SQL-aided table understanding. CoQ adopts natural-language-style representations of table schemas to abstract away structural noise and enhance understanding. It employs a clause-by-clause SQL generation strategy to improve query quality and introduces a hybrid reasoning division that separates SQL-based mechanical reasoning from LLM-based logical inference, thereby reducing reliance on execution outcomes. Extensive experiments across four models and five widely used benchmarks demonstrate that CoQ achieves substantial accuracy improvements and significantly lowers invalid SQL rates compared to prior generic LLM-based, SQL-aided, and hybrid baselines, confirming its superior effectiveness in table understanding. The code is available at https://github.com/SongyuanSui/ChainofQuery.
@@
Songyuan Sui1, Hongyi Liu1, Serena Liu1, Li Li2, Soo-Hyun Choi3, Rui Chen2, Xia Hu1 1Rice University {Songyuan.Sui, Hongyi.Liu, Serena.Liu, Xia.Hu}@rice.edu 2Samsung Electronics America 3Warner Bros. Discovery {li.li1, rui.chen1}@samsung.com soohyun.choi@wbd.com
@@ 1 Introduction
Large Language Models (LLMs) have shown remarkable performance across a variety of natural language processing (NLP) tasks (yang2023harnessingpowerllmspractice). However, they still struggle with understanding tabular data (sui2023table). This challenge stems from two main factors. First, the structure of tabular data differs significantly from that of plain text. It organizes information via rows and columns, introducing hierarchical relationships, positional dependencies, and implicit semantics that are challenging for language models to capture. Second, table-based tasks often involve multi-step, structured reasoning operations, such as aggregation, comparison, and arithmetic computation (lu2025large). They go beyond surface-level language understanding.
To adapt LLMs for table understanding, prior work has explored many approaches. Many studies (Herzig_2020; zhang2023generative; zhang2023tablellama; li2024tablegpt; he2025tablelora) fine-tuned LLMs on table-specific datasets, but these methods treat the task as a single-turn generation problem, which results in shallow reasoning paths and overlooks intermediate data. Recent studies (wang2024chaintab; ji2024treetab; zhou2025efficientmultiagentcollaborationtool) built agent-based pipelines with table manipulation tools to conduct multi-step reasoning. Nevertheless, these tools rely on human-defined programs whose functionalities are restricted and cannot perform complex or adaptive operations beyond their predefined scope.
@@
We argue that incorporating Structured Query Language (SQL) provides a more principled approach for understanding tabular data. SQL is inherently designed to access, filter, and aggregate information from relational tables, making it well-aligned with the needs of table understanding. Although existing work on LLM-based SQL generation has made some progress in table understanding (ye2023dater; kong2024opentab; huang2024queryagentreliableefficientreasoning; abhyankar2025hstarllmdrivenhybridsqltext), its performance remains suboptimal. We identify three key challenges in SQL-aided table understanding with LLMs. 1. Existing approaches typically feed the table and task instructions into an LLM, assuming the LLM can understand the table in order to produce meaningful SQL. However, if the LLM already struggles with table understanding, it is unlikely to generate accurate SQL in the first place. This workflow thus inherits LLMs’ structural inability, as effective SQL generation still requires a solid grasp of the table’s layout and semantics—something LLMs inherently lack. 2. Supporting complex, multi-step reasoning often requires SQL queries with multiple clauses and nested structures. Such queries are difficult to construct correctly in a single pass, and even small errors can cascade into query failure. 3. Existing methods depend heavily on the correctness of SQL execution, as they directly use SQL results as final answers. This further leads LLMs to generate overly complicated queries, thereby reducing overall reliability.
@@
To tackle these challenges, we propose Chain-of-Query (CoQ), a novel multi-agent framework for SQL-aided table understanding. To address the first challenge, CoQ replaces raw table inputs with natural-language representations of table schemas. Our insight is that SQL generation depends primarily on high-level schemas, which can be easily and accurately expressed in natural language, bypassing the need for LLMs to interpret complex tabular structures. This abstraction allows LLMs to focus on semantics without being hindered by noisy or irregular layouts. To address the second challenge, we propose a Clause-by-Clause SQL Generation Strategy, which incrementally constructs SQL queries one clause at a time, forming a chain of progressively refined queries. To solve the third challenge, we design a Hybrid Reasoning Division Strategy that separates mechanical reasoning (executed via SQL) from logical reasoning (delegated to LLMs), treating SQL outputs as intermediate steps rather than final answers. In addition, our Parallel Decomposition enhances robustness by splitting complex questions into parallelizable sub-questions, avoiding the inter-step dependencies of traditional sequential reasoning. Together, CoQ offers a novel design that directly addresses these challenges, establishing an effective new paradigm for table understanding.
@@
To comprehensively evaluate CoQ, we conduct extensive experiments with both closed-source models (GPT-3.5, GPT-4.1) and open-source models (LLaMA 2, DeepSeek-V3) on five representative benchmarks: WikiTQ, TabFact, FeTaQA, IM-TQA, and Open-WikiTable. CoQ is compared with strong baselines across four categories: generic LLM-based, SQL-aided, hybrid, and general structured data reasoning. Results show that CoQ consistently outperforms state-of-the-art (SOTA) methods; for example, on WikiTQ it improves accuracy from 61.11% to 74.77% and reduces invalid-SQL rates from 9.48% to 3.34% with GPT-3.5.
@@
  • Extensive experiments show that Chain-of-Query consistently outperforms previous baselines, achieving substantial improvements across diverse settings.
@@ 2 Related Work
We review related work in three areas: LLM-based table understanding, Text-to-SQL, and hybrid methods for table reasoning. Additional discussions on multi-agent approaches for table understanding and general structured data reasoning with LLMs can be found in Appendix I.
LLM-Based Table Understanding. Since the structure of tables (e.g., non-sequential cell order) differs significantly from plain text typically used in pre-training (raffel2023exploringlimitstransferlearning), several studies (zhang2023generative; zhang2023tablellama; li2024tablegpt; zhuang2024structlmbuildinggeneralistmodels; he2025tablelora) fine-tuned LLMs to adapt to tabular data. While effective on specific datasets, these methods depend on manually crafted instructions and expensive training, limiting scalability. More recent prompt-based approaches (sui2023table; sui2023tap4llm; chen2024tablerag) exploit LLMs’ general reasoning ability, offering better generalization across tasks. However, these methods typically adopt single-turn prompting strategies, failing to support multi-hop reasoning.
Text-to-SQL. Generating SQL queries from natural language is a long-standing challenge. Early methods primarily relied on fine-tuning LLMs on annotated SQL datasets (pourreza2024dtssqldecomposedtexttosqlsmall; li2024codesbuildingopensourcelanguage). To reduce training cost and improve compositional reasoning, tai2023exploringchainofthoughtstyleprompting and pourreza2023dinsqldecomposedincontextlearning explored Chain-of-Thought prompting in Text-to-SQL. OpenTab (kong2024opentab) integrates LLM-based SQL generation into a tabular data RAG system. Recently, MAC-SQL (wang2025macsql) and MAG-SQL (xie2024mag) build multi-agent frameworks that refine generated queries to improve SQL quality. Although these methods perform well on benchmarks such as Spider (yu2019spiderlargescalehumanlabeleddataset) and BIRD (li2023llmservedatabaseinterface), they all tend to generate fully grounded SQL queries that directly return the final answer, making query construction difficult. Moreover, LLMs tend to produce overly complex queries in an attempt to handle questions holistically (shen2025studyincontextlearningbasedtexttosqlerrors), reducing reliability.
Hybrid Methods for Table Understanding. To address the shortcomings of LLMs in structured reasoning, hybrid approaches integrate SQL with LLMs. ReAcTable (zhang2023reactableenhancingreacttable) generates multiple SQL candidates for intermediate information retrieval but at the cost of increased computation and limited structural understanding. SynTQA (zhang2024syntqasynergistictablebasedquestion) generates both Text-to-SQL and LLM answers but only selects one result, discarding complementary information. Recent studies (nahid2024tabsqlifyenhancingreasoningcapabilities; cao2025tablemasterrecipeadvancetable) use SQL to select key columns and rows before invoking LLMs. While effective for data reduction, they apply SQL only at a basic level, overlooking its potential for complex table manipulations. H-STAR (abhyankar2025hstarllmdrivenhybridsqltext) further employs complex SQL-based symbolic reasoning, but its reliance on single-shot SQL generation at each stage makes it vulnerable to execution failures.
@@ 3.3 Semantic Splitter: Schema Abstraction and Query Decomposition
The Semantic Splitter serves as the entry point of the entire agentic workflow. This module is designed to optimize the input for SQL generation by addressing two key challenges: (1) LLMs’ limited ability to interpret complex table structures, which often leads to inaccurate SQL generation; and (2) interference between sub-questions within a single complex query, where independent sub-questions could be handled separately to avoid entanglement. To address these challenges, we introduce a natural-language-style representation of table schemas for query generation, and incorporate a parallel decomposition mechanism to identify and isolate separable sub-queries for clean execution. The following subsections detail this process. An illustrative example is provided in Appendix B.1.
@@ 3.3.1 Natural-Language-Style Table Schema
To address these limitations in handling raw and irregular table structures, we construct a concise natural-language-style schema as a surrogate for full-table input. It consists of three components: table name, headers, and value examples. The table name employs concise keywords to summarize the main content and purpose of the table, providing essential context for the LLM to understand the table. The table headers comprise multiple (column name, column value type) pairs. These pairs not only provide a natural-language abstraction of the table structure, but also serve as a reference for value formatting in downstream SQL generation. The value examples are a few sampled rows of the table paired with their corresponding column names. We adopt "PIPE" format (i.e., a pipe-delimited string with "|" separators) to encode the value examples, as studies found it to improve LLM performance on table tasks (sui2023table; wang2024chaintab).
By replacing the full table with our designed schema, LLMs can generate accurate SQL without being distracted by complex layouts or unnecessary content. This approach also improves scalability by avoiding full-table imports, thereby reducing input length and simplifying reasoning.
@@ 3.3.2 Parallel Decomposition
The second component of the Semantic Splitter is a Parallel Decomposer, which focuses on separating input questions into independent sub-questions. Unlike traditional sequential decomposition strategies, where each sub-question depends on the previous answer and becomes fragile due to error propagation, our approach focuses on questions without interdependencies. We defer the handling of sequential dependencies to our Clause-by-Clause Generator in Section 3.4.2.
Many complex tabular questions involve comparing or aggregating information from semantically disjoint table regions, which do not require sequential reasoning. To address this, the Parallel Decomposer splits such questions into independent sub-questions, each targeting a localized region of the table. This removes inter-step dependencies and enables concurrent processing. Each sub-query then operates over a narrow, self-contained scope without interference from others, simplifying SQL generation.
@@ 3.4 Clause-wise SQL Generation for Precision and Robustness
After outlining our input transformation and preprocessing steps, we now focus on the next key component—generating accurate SQL with LLMs to support effective table understanding. The SQL Query Generator constructs executable SQL queries to retrieve intermediate information relevant to the task and support mechanical reasoning. As shown in Figure 2, an initial query is generated by an LLM, then validated before being passed to the Dynamic Planner, which determines whether further clauses should be appended to improve precision. The following subsection details the generation process.
@@ 3.4.1 SQL Query Generation with Validation
Given the schema and sub-question, an SQL query is first generated by an LLM to retrieve relevant information. The query is validated by directly executing it via an SQL executor; if it fails, the error message is passed to a secondary LLM for correction. The revised query is re-executed and, if successful, forwarded to the Dynamic Planner. To enhance query quality, we introduce a clause-by-clause generation strategy described below.
@@ 3.4.2 Clause-by-Clause SQL Generation
To improve the reliability and controllability of SQL generation, we propose a Clause-by-Clause Strategy that incrementally builds queries by isolating the construction of each clause. The generator first produces a basic SELECT-FROM query selecting columns relevant to the sub-question. Without condition filtering, however, this query may return a large number of rows, posing scalability challenges. To mitigate this, we sample a few representative rows (similar to the natural-language-style table schema construction) as the Planner only needs to observe data patterns rather than precise values.
Subsequent clauses (e.g., WHERE, GROUP BY) are appended one at a time, under the guidance of the Dynamic Planner. By decomposing SQL construction into discrete clauses, each step focuses solely on its own logic and syntax, independent of the full query context. Each intermediate query is validated before proceeding. If execution fails, the error can be traced to the most recently added clause, allowing for more targeted correction. If correction fails, the agent reverts to the last validated query, ensuring robustness. As shown in Figure 2, if a newly generated clause introduces an error to the query (SQL2) and cannot be resolved, the agent uses the last valid query (SQL1) to ensure executable SQL. This incremental process forms a chain of increasingly precise queries, where each step builds on a validated state, preventing error propagation and ensuring stability. The list of clauses is provided in Appendix C.